OpenStudy (anonymous):

What is the solution to the system of equations that contains −3x + y = 3 and 2x − y = −1? last question..

OpenStudy (imtiaz7):

these two are linear equations and can be solved by substitution method easily

OpenStudy (imtiaz7):

from 1st eq. -3x+y=3 y=3x+3 (we get the value of y) now put it in 2nd eq. 2x-y=-1 2x-(3x+3)=-1 2x-3x-3=-1 -x-3=-1 x=-3+1 x=-2 now we get x and for y put x's value in any one eq. take eq. 1: -3x+y=3 -3(-2)+y=3 6+y=3 y=3-6 y=-3 Hence we get x=-2 and y=-3

OpenStudy (anonymous):

@Imtiaz7 nevermind i put the wrong answer choices

OpenStudy (imtiaz7):

uff you confused me so much but it's ok you accepted your mistake :)

OpenStudy (imtiaz7):

so it correct?

OpenStudy (imtiaz7):

@kimm210

OpenStudy (anonymous):

@Imtiaz7 yes it was correct , got a 100 :)

OpenStudy (imtiaz7):

ok :)